# Part I: General framework, application to Hartree-Fock and DFT

### v1

@@ 4.2 Analysis at the UHF level of theory
Let us now consider the UHF approximation, for which the underlying manifold is $\mathcal{M}={\rm Gr}_{\mathbb{C}}(N_{\rm p},N_{\rm b})\times{\rm Gr}_{\mathbb{C}}(N_{\rm p},N_{\rm b})$. The UHF approximation of $\widehat{H}_{\lambda}$ gives rise to a parametrized family of energy functionals of the form
$E^{\rm UHF}(\lambda,\gamma_{\alpha},\gamma_{\beta})=$ ${\rm Tr\,}(h\gamma_{\alpha})+{\rm Tr\,}(h\gamma_{\beta})+\frac{\lambda}{2}{\rm Tr\,}(J(\gamma_{\alpha}+\gamma_{\beta})(\gamma_{\alpha}+\gamma_{\beta}))$
$\quad-\frac{\lambda}{2}{\rm Tr\,}(K(\gamma_{\alpha})\gamma_{\alpha})-\frac{\lambda}{2}{\rm Tr\,}(K(\gamma_{\beta})\gamma_{\beta}),$ (43)
@@
CP-UHF. The critical points of $\mathcal{M}\ni(\gamma_{\alpha},\gamma_{\beta})\mapsto\mathcal{E}^{\rm UHF}(0,\gamma_{\alpha},\gamma_{\beta})={\rm Tr\,}(h\gamma_{\alpha})+{\rm Tr\,}(h\gamma_{\beta})\in{\mathbb{R}}$ are the $(\gamma_{\alpha\star},\gamma_{\beta\star})\in\mathcal{M}$ satisfying $[h,\gamma_{\alpha\star}]=[h,\gamma_{\beta\star}]=0$. Under the nondegeneracy assumption (H), these critical points are isolated and there are exactly $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ of them. The ground state is
@@
for $I_{\sigma}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\sigma}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I|=|A|\geq 1$. It follows from a simple application of the analytic implicit function theorem that under the nondegeneracy assumption (H), there exist $\lambda_{\rm c}>0$ and $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ real-analytic functions
$\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\bullet),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\bullet)\right):(-\lambda_{\rm c},\lambda_{\rm c})\to\mathcal{M},$
$I_{\alpha},I_{\beta}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\alpha},A_{\beta}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I_{\alpha}|=|A_{\alpha}|$, $|I_{\beta}|=|A_{\beta}|$, such that for any $\lambda\in(-\lambda_{\rm c},\lambda_{\rm c})$, the function $\mathcal{M}\ni(\gamma_{\alpha},\gamma_{\beta})\mapsto\mathcal{E}(\lambda,\gamma_{\alpha},\gamma_{\beta})\in{\mathbb{R}}$ has exactly $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ critical points, all isolated, nondegenerate, and given by $\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\lambda),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\lambda)\right)$. Note that, by symmetry, $\gamma_{\alpha,[\emptyset,\emptyset]}(\lambda)=\gamma_{\beta,[\emptyset,\emptyset]}(\lambda)=:\gamma_{[\emptyset,\emptyset]}(\lambda)$, i.e., the UHF ground state coincides with the RHF ground state.
For $\lambda\in(-\lambda_{\rm c},\lambda_{\rm c})$, the excitation energies obtained by CP-UHF are thus the $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}-1$ positive numbers
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}(\lambda):=E^{\rm UHF}\left(\lambda,\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\lambda),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\lambda)\right)-E^{\rm UHF}\left(\lambda,\gamma_{[\emptyset,\emptyset]}(\lambda),\gamma_{[\emptyset,\emptyset]}(\lambda)\right)$ (46)
with $I_{\sigma}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\sigma}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I_{\sigma}|=|A_{\sigma}|$, $|I_{\alpha}|+|I_{\beta}|=|A_{\alpha}|+|A_{\beta}|\geq 1$. In addition, the functions $\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}:(-\lambda_{\rm c},\lambda_{\rm c})\to{\mathbb{R}}$ are real-analytic and it holds
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}(\lambda)=\sum_{k=0}^{+\infty}\lambda^{k}\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF},(k)},$
@@
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF},(1)}$ $=\frac{\partial E^{\rm UHF}}{\partial\lambda}\left(0,\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}^{(0)},\gamma_{\beta,[I_{\beta},A_{\beta}]}^{(0)}\right)-\frac{\partial E^{\rm UHF}}{\partial\lambda}\left(0,\gamma^{(0)}_{[\emptyset,\emptyset]},\gamma^{(0)}_{[\emptyset,\emptyset]}\right).$
@@
$=$ $\sum_{i=1}^{N_{\rm p}-1}((N_{\rm p}+1)\sigma,(N_{\rm p}+1)\sigma||i\sigma,i\sigma)+\sum_{i=1}^{N_{\rm p}}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta|i\alpha,i\alpha)$
$-\sum_{i=1}^{N_{\rm p}-1}(N_{\rm p}\sigma,N_{\rm p}\sigma||i\sigma,i\sigma)-\sum_{i=1}^{N_{\rm p}}(N_{\rm p}\beta,N_{\rm p}\beta|i\alpha,i\alpha)$ (47)
@@
LR-UHF. Let us now turn to the LR-UHF approximation of excitation energies. First, as the complex dimension of the manifold $\mathcal{M}$ is $2N_{\rm p}(N_{\rm b}-N_{\rm p})$, there are at most $2N_{\rm p}(N_{\rm b}-N_{\rm p})$ of them. For $\lambda=0$, they can be interpreted as single excitations:
@@
Using again the analytic implicit functional theorem, it can be shown that for $\lambda$ small enough, the excitation energies computed by linear response are given by real-analytic functions $(-\lambda_{\rm c},\lambda_{\rm c})\ni\lambda\mapsto\omega_{[i,a]}^{{\rm LR-UHF}}(\lambda)\in{\mathbb{R}}$ such that $\omega_{[i,a]}^{{\rm LR-UHF}}(0)=\omega_{[i,a]}^{{\rm LR-UHF},(0)}$. In particular, the first excited state corresponds to the HOMO-LUMO transition and the first-order variation is given by
$\omega^{\rm LR-UHF,(1)}_{\rm HOMO-LUMO}:=\omega^{\rm LR-UHF,(1)}_{[N_{\rm p},N_{\rm p}+1]}$
$=$ $\sum_{j=1}^{N_{\rm p}-1}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta||j\beta,j\beta)-\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p}\beta,N_{\rm p}\beta||j\beta,j\beta)$
$+\sum_{j=1}^{N_{\rm p}}((N_{\rm p}+1)\beta,(N_{\rm p}+1)\beta|j\alpha,j\alpha)-\sum_{j=1}^{N_{\rm p}}(N_{\rm p}\beta,N_{\rm p}\beta|j\alpha,j\alpha)$
$-(N_{\rm p}\beta,(N_{\rm p}+1)\beta|(N_{\rm p}+1)\alpha,N_{\rm p}\alpha).$ (48)
Notice that this expression coincides with the one found in the FCI case, in Eq. (42). Therefore, in the non-interacting limit, FCI and UHF-LR are equivalent in computing single excitation energies, while they differ in the interacting case.
@@ 5 Numerical simulations
We perform proof-of-concept numerical simulations to support our theoretical findings. In Section 5.1, we compare the CP and LR formalisms in UHF theory for the H2, H4, and H2O molecules, with a varying coupling parameter $\lambda$. The exact FCI theory is taken as reference. In particular, we check the analytical first-order derivatives of the excitation energies at $\lambda=0$ given by Eqs. (47) and (48) vs the numerical ones obtained by local linear fits. In Section 5.2, we demonstrate that multiple critical points can arise even for the H4 molecule (rectangular geometry) at the UHF/STO-3G level of theory. The energies of the critical points and their overlaps with the exact FCI excited states are shown as functions of $\lambda$. We note that in this section, only the index-1 saddle points (namely, the critical points where the Riemannian Hessian has exactly one negative eigenvalue) are considered in the CP formalism and they are computed numerically by using a constrained gentlest ascent dynamics, which is described in detail in another ongoing work. The approximate excitation energies in the LR formalism are found by solving Casida’s equations 23, 24, 25 using the open-source package PySCF 56.
@@ 5.1 Comparison of CP and LR formalisms within the UHF theory
In this Section, numerical experiments performed in the UHF framework are shown. Calculations have been performed with the minimal basis set STO-3G 57 and with the 3-21G basis 58, for the systems H2, H4 in a linear and a rectangular geometry, and H2O, with $\lambda$ in the range $[0,1]$. For H2, a bond length of 1 a.u. was considered; for the H4 molecule, the linear geometry considers a bond length of 0.875 Å (Ref. 59); for the rectangular geometry, sides of 1 Å and 2 Å are considered. For the water molecule, the coordinates [O (0.00000000 0.00000000 0.11993333) a.u.], [H ($00$1.43497461 -0.95171452) a.u.], [H (0.00000000 1.43497461 -0.95171452) a.u.] were taken. We compare the first-order excitation energies of the HOMO-LUMO transition, obtained both with the LR and CP approaches. We note that for the CP results in this subsection, only the lowest excited states at $\lambda=0$, which correspond to the HOMO-LUMO excitations, have been tracked across the range of $\lambda$. When $\lambda$ increases, multiple critical points (with the same index) may emerge; see the discussions in Section 5.2.
@@
Table 1: First-order excitation energies for the HOMO-LUMO excitations of the molecules H2, H4 in linear and rectangular geometry, and H2O, obtained in the non-interacting regime of $\lambda\xrightarrow[]{}0$. For both CP and LR, the results obtained by the analytic expressions Eqs. (47) and (48), denoted here by $\omega_{\rm anal}^{{\rm CP},(1)}$ and $\omega_{\rm anal}^{{\rm LR},(1)}$, and numerically using a five-point scheme linear fit, denoted here by $\omega_{\rm num}^{{\rm CP},(1)}$ and $\omega_{\rm num}^{{\rm LR},(1)}$, are listed. The difference $\Delta\omega_{\rm anal}^{(1)}=\omega_{\rm anal}^{{\rm CP},(1)}-\omega_{\rm anal}^{{\rm LR},(1)}$ between the two approaches is also shown.
Molecule (Basis) $\omega_{\rm num}^{{\rm CP},(1)}$ $\omega_{\rm anal}^{{\rm CP},(1)}$ $\omega_{\rm num}^{{\rm LR},(1)}$ $\omega_{\rm anal}^{{\rm LR},(1)}$ $\Delta\omega_{\rm anal}^{(1)}$
H2 (STO-3G) -0.012580 -0.012586 -0.182839 -0.182827 0.170241
H2 (3-21G) -0.330913 -0.330929 -0.422845 -0.422889 0.091960
H4 lin. (STO-3G) -0.107158 -0.107347 -0.245692 -0.245812 0.138465
H4 lin. (3-21G) -0.207375 -0.207529 -0.332364 -0.332485 0.124956
H4 rect. (STO-3G) 0.052118 0.052118 -0.040455 -0.040246 0.092364
H4 rect.(3-21G) -0.061798 -0.061775 -0.133340 -0.133204 0.071429
H2O (STO-3G) -3.494042 -3.494149 -3.497100 -3.497208 0.003059
@@
The results displayed in Table 1 show for all examples a good agreement between the analytical derivatives obtained in Section 4 and linear fits of these quantities using the five points $(j\cdot\delta\lambda)_{0\leq j\leq 4}$, with $\delta\lambda=10^{-4}$ chosen small enough to be compatible with the limit $\lambda\xrightarrow[]{}0$, but not too small to avoid numerical instabilities. Furthermore, we can observe that, for the examples considered here, the difference calculated as $\omega_{\mathrm{anal}}^{{\rm CP},(1)}-\omega_{\mathrm{anal}}^{{\rm LR},(1)}$ is always positive, showing that the LR result is always smaller than the CP one for the explored systems.
@@
Figure 2: Dependence on the $\lambda$-parameter of the excitation energy, calculated on the H2 molecule, with the 3-21G basis set.
@@
Figure 3: Dependence on the $\lambda$-parameter of the excitation energy, calculated on the H4 molecule, in the linear (a) and rectangular (b) geometries, with the 3-21G basis set.
@@
The two approaches CP and LR, at the UHF level of theory can be compared with the exact FCI results. The results are shown in Figs. 2-4, for the molecules H2 and H4 in the linear and rectangular geometry, with the 3-21G basis set 58.
@@
Figure 4: Dependence on the $\lambda$-parameter of the excitation energy, calculated on the H2O, with the STO-3G basis set. In Fig. 4(b), a zoom on the differences between the different approaches in the region around $\lambda=1$ is plotted.
@@
Figure 5: Dependence on the $\lambda$-parameter of the excitation energy, calculated on H2O. In Fig. 5(a) the 6-31G basis set is used, while in Fig. 5(b), the cc-pVDZ basis set is used. At $\lambda=0.2$ and $\lambda=0.6$, the TD-HF algorithm in PySCF was not able to converge due to high linear dependencies.
@@
For all cases, the LR results are closer to the FCI ones. The plots also show that, in the non-interacting limit ($\lambda=0$), the LR formalism coincides with FCI as expected from Eq. (48), while the two approaches differ in the interacting case. In the water example with the STO-3G (Fig. 4) 57, 6-31G 60, 61, and cc-pVDZ 62 basis sets (Fig. 5), the three methods show small discrepancies across the range until $\lambda$ approaches 1.

### v2

@@ 4.2 Analysis at the UHF level of theory
Let us now consider the UHF approximation, for which the underlying manifold is $\mathcal{M}={\rm Gr}_{\mathbb{C}}(N_{\rm p},N_{\rm b})\times{\rm Gr}_{\mathbb{C}}(N_{\rm p},N_{\rm b})$. The UHF approximation of $\widehat{H}_{\eta}$ gives rise to a parametrized family of energy functionals of the form
$E^{\rm UHF}_{\eta}(\gamma_{\alpha},\gamma_{\beta})=$ ${\rm Tr\,}(h\gamma_{\alpha})+{\rm Tr\,}(h\gamma_{\beta})+\frac{\eta}{2}{\rm Tr\,}(J(\gamma_{\alpha}+\gamma_{\beta})(\gamma_{\alpha}+\gamma_{\beta}))$
$\quad-\frac{\eta}{2}{\rm Tr\,}(K(\gamma_{\alpha})\gamma_{\alpha})-\frac{\eta}{2}{\rm Tr\,}(K(\gamma_{\beta})\gamma_{\beta}),$ (46)
@@
CP-UHF. The critical points of $\mathcal{M}\ni(\gamma_{\alpha},\gamma_{\beta})\mapsto\mathcal{E}^{\rm UHF}_{0}(\gamma_{\alpha},\gamma_{\beta})={\rm Tr\,}(h\gamma_{\alpha})+{\rm Tr\,}(h\gamma_{\beta})\in{\mathbb{R}}$ are the $(\gamma_{\alpha\star},\gamma_{\beta\star})\in\mathcal{M}$ satisfying $h,\gamma_{\alpha\star}]=[h,\gamma_{\beta\star}]=0$. Under the nondegeneracy assumption (H), these critical points are isolated and there are exactly $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ of them. The ground state is
@@
for $I_{\sigma}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\sigma}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I|=|A|\geq 1$. It follows from a simple application of the analytic implicit function theorem that under the nondegeneracy assumption (H), there exist $\eta_{\rm c}>0$ and $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ real-analytic functions
$\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\bullet),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\bullet)\right):(-\eta_{\rm c},\eta_{\rm c})\to\mathcal{M},$
$I_{\alpha},I_{\beta}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\alpha},A_{\beta}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I_{\alpha}|=|A_{\alpha}|$, $|I_{\beta}|=|A_{\beta}|$, such that for any $\eta\in(-\eta_{\rm c},\eta_{\rm c})$, the function $\mathcal{M}\ni(\gamma_{\alpha},\gamma_{\beta})\mapsto\mathcal{E}_{\eta}(\gamma_{\alpha},\gamma_{\beta})\in{\mathbb{R}}$ has exactly $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}$ critical points, all isolated, nondegenerate, and given by $\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\eta),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\eta)\right)$. Note that, by symmetry, $\gamma_{\alpha,[\emptyset,\emptyset]}(\eta)=\gamma_{\beta,[\emptyset,\emptyset]}(\eta)=:\gamma_{[\emptyset,\emptyset]}(\eta)$, i.e., the UHF ground state coincides with the RHF ground state.
For $\eta\in(-\eta_{\rm c},\eta_{\rm c})$, the excitation energies obtained by CP-UHF are thus the $\binom{N_{\rm b}}{N_{\rm p}}\times\binom{N_{\rm b}}{N_{\rm p}}-1$ positive numbers
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}(\eta):=E^{\rm UHF}_{\eta}\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}(\eta),\gamma_{\beta,[I_{\beta},A_{\beta}]}(\eta)\right)-E^{\rm UHF}_{\eta}\left(\gamma_{[\emptyset,\emptyset]}(\eta),\gamma_{[\emptyset,\emptyset]}(\eta)\right)$ (49)
with $I_{\sigma}\subset\{1,\ldots,N_{\rm p}\}$, $A_{\sigma}\subset\{N_{\rm p}+1,\ldots,N_{\rm b}\}$, $|I_{\sigma}|=|A_{\sigma}|$, $|I_{\alpha}|+|I_{\beta}|=|A_{\alpha}|+|A_{\beta}|\geq 1$. In addition, the functions $\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}:(-\eta_{\rm c},\eta_{\rm c})\to{\mathbb{R}}$ are real-analytic and it holds
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF}}(\eta)=\sum_{k=0}^{+\infty}\eta^{k}\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF},(k)},$
@@
$\omega_{[I_{\alpha},A_{\alpha}],[I_{\beta},A_{\beta}]}^{{\rm CP-UHF},(1)}$ $=\left.\frac{\partial E^{\rm UHF}_{\eta}}{\partial\eta}\right|_{\eta=0}\left(\gamma_{\alpha,[I_{\alpha},A_{\alpha}]}^{(0)},\gamma_{\beta,[I_{\beta},A_{\beta}]}^{(0)}\right)-\left.\frac{\partial E^{\rm UHF}_{\eta}}{\partial\eta}\right|_{\eta=0}\left(\gamma^{(0)}_{[\emptyset,\emptyset]},\gamma^{(0)}_{[\emptyset,\emptyset]}\right).$
@@
$=$ $\sum_{i=1}^{N_{\rm p}-1}(N_{\rm p}+1,N_{\rm p}+1||i,i)_{\sigma\sigma}+\sum_{i=1}^{N_{\rm p}}(N_{\rm p}+1,N_{\rm p}+1|i,i)_{\beta\alpha}$
$-\sum_{i=1}^{N_{\rm p}-1}(N_{\rm p},N_{\rm p}||i,i)_{\sigma\sigma}-\sum_{i=1}^{N_{\rm p}}(N_{\rm p},N_{\rm p}|i,i)_{\beta\alpha}$ (50)
@@
LR-UHF. Let us now turn to the LR-UHF approximation of excitation energies. As the complex dimension of the manifold $\mathcal{M}$ is $2N_{\rm p}(N_{\rm b}-N_{\rm p})$, there are at most $2N_{\rm p}(N_{\rm b}-N_{\rm p})$ of them. For $\eta=0$, they can be interpreted as single excitations:
@@
In analogy with the FCI case, there are $N_{\rm p}(N_{\rm b}-N_{\rm p})$ two-fold degenerate excitation energies at $\eta=0$ due to the symmetry resulting from swapping spin-up and spin-down electrons. Likewise, the first-order variation of the lowest excitation energy is given by the triplet excited state as
$\omega_{\rm HOMO-LUMO}^{\rm LR-UHF,(1)}=$ $\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p}+1,N_{\rm p}+1||j,j)_{\beta\beta}-\sum_{j=1}^{N_{\rm p}-1}(N_{\rm p},N_{\rm p}||j,j)_{\beta\beta}$
$+\sum_{j=1}^{N_{\rm p}}(N_{\rm p}+1,N_{\rm p}+1|j,j)_{\beta\alpha}-\sum_{j=1}^{N_{\rm p}}(N_{\rm p},N_{\rm p}|j,j)_{\beta\alpha}$
$-(N_{\rm p},N_{\rm p}+1|N_{\rm p}+1,N_{\rm p})_{\beta\alpha}.$ (51)
Notice that this expression coincides with the one found in the FCI case, in Eq. (45). At first order, FCI thus coincides with LR in the UHF setting, both of which allow symmetry-breaking. However, when considering the restricted HF (RHF) case, for closed-shell systems, the expressions change and spin degeneracy is no longer an issue. For reference, we can construct a FCI in theory built only from Slater determinants with doubly occupied orbitals and will temporarily refer to it as restricted FCI here. In this case, restricted FCI can be shown to coincide with CP-RHF in the weakly interacting regime, but different from LR-RHF. Lastly, if we consider the general HF (GHF) in the spinless setting, it can be shown that all the three methods, FCI, LR-GHF, and CP-GHF coincide at $\eta=0$ up to the first order.
@@ 5 Numerical simulations
We perform proof-of-concept numerical simulations to support our theoretical findings. In Section 5.1, we compare the CP and LR formalisms in UHF theory for the H2, H4, and H2O molecules, with a varying coupling parameter $\eta$. The exact FCI theory is taken as reference. In particular, we check the analytical first-order derivatives of the excitation energies at $\eta=0$ given by Eqs. (50) and (51) vs the numerical ones obtained by finite differences. In Section 5.2, we demonstrate that multiple critical points can arise even for the H4 molecule (rectangular geometry) at the UHF/STO-3G level of theory. The energies of the critical points and their overlaps with the exact FCI excited states are shown as functions of $\eta$. We note that in this section, only the index-1 saddle points (namely, the critical points where the Riemannian Hessian has exactly one negative eigenvalue) are considered in the CP formalism and they are computed numerically by using a constrained gentlest ascent dynamics, which is described in detail in a paper in preparation. The approximate excitation energies in the LR formalism are found by solving Casida’s equations [23, 24, 25] using the open-source package PySCF.[70]
@@ 5.1 Comparison of CP and LR formalisms within the UHF theory
In this subsection, numerical experiments performed in the UHF framework are shown. Calculations have been performed with the minimal basis set STO-3G[71] and with the 3-21G basis,[72] for the systems H2, H4 in a linear and a rectangular geometry, and H2O, with $\eta$ in the range $[0,1]$. For H2, a bond length of 1 a.u. was considered; for the H4 molecule, the linear geometry considers a bond length of 0.875 Å (Ref. [73]); for the rectangular geometry, sides of 1 Å and 2 Å are considered. For the water molecule, the coordinates [O (0.00000000 0.00000000 0.11993333) a.u.], [H (0.00000000 -1.43497461 -0.95171452) a.u.], [H (0.00000000 1.43497461 -0.95171452) a.u.] were taken. We compare the first-order variation of the lowest excitation energies, obtained with both the LR and CP approaches. We note that for the CP results in this subsection, only the lowest excited states at $\eta=0$, which correspond to the HOMO-LUMO excitations, have been tracked across the range of $\eta$. When $\eta$ increases, multiple critical points (with the same index) may emerge; see the discussions in Section 5.2.
@@
Table 2: First-order variations for the lowest excitation energies of the molecules H2, H4 in linear and rectangular geometry, and H2O, obtained in the non-interacting regime of $\eta\xrightarrow[]{}0$. For both CP and LR, the results obtained by the analytic expressions Eqs. (50) and (51), denoted here by $\omega_{\rm anal}^{{\rm CP},(1)}$ and $\omega_{\rm anal}^{{\rm LR},(1)}$, and numerically using a finite-difference scheme, denoted here by $\omega_{\rm num}^{{\rm CP},(1)}$ and $\omega_{\rm num}^{{\rm LR},(1)}$, are listed. The difference $\Delta\omega_{\rm anal}^{(1)}=\omega_{\rm anal}^{{\rm CP},(1)}-\omega_{\rm anal}^{{\rm LR},(1)}$ between the two approaches is also shown.
Molecule (Basis) $\omega_{\rm num}^{{\rm FCI},(1)}$ $\omega_{\rm anal}^{{\rm FCI},(1)}$ $\omega_{\rm num}^{{\rm CP},(1)}$ $\omega_{\rm anal}^{{\rm CP},(1)}$ $\omega_{\rm num}^{{\rm LR},(1)}$ $\omega_{\rm anal}^{{\rm LR},(1)}$ $\Delta\omega_{\rm anal}^{(1)}$
H2 (STO-3G) -0.18283 -0.18283 -0.01259 -0.01259 -0.18283 -0.18283 0.17024
H2 (3-21G) -0.42288 -0.42289 -0.33093 -0.33093 -0.42288 -0.42289 0.09196
H4 lin. (STO-3G) -0.24580 -0.24581 -0.10733 -0.10735 -0.24580 -0.24581 0.13847
H4 lin. (3-21G) -0.33247 -0.33249 -0.20751 -0.20753 -0.33247 -0.33249 0.12496
H4 rect. (STO-3G) -0.04026 -0.04025 0.05212 0.05212 -0.04027 -0.04025 0.09236
H4 rect.(3-21G) -0.13321 -0.13320 -0.06178 -0.06178 -0.13322 -0.13320 0.07143
H2O (STO-3G) -3.49415 -3.49721 -3.49414 -3.49415 -3.49720 -3.49721 0.00306
The results displayed in Table 2 show for all examples a good agreement between the analytical derivatives obtained in Section 4 and the numerical results obtained through a finite-difference scheme. The latter are calculated by
$\omega_{\rm num}^{(1)}=\frac{\omega(\delta\eta)-\omega(0)}{\delta\eta}.$
@@
In these results, $\delta\eta$ has been chosen to be $10^{-4}$. Furthermore, we can observe that, for the examples considered here, the difference calculated as $\omega_{\mathrm{anal}}^{{\rm CP},(1)}-\omega_{\mathrm{anal}}^{{\rm LR},(1)}$ is always positive, showing that the LR result is always smaller than the CP one for the explored systems.
@@
Figure 3: Dependence on the $\eta$-parameter of the excitation energy, calculated on the H2 molecule, with the 3-21G basis set.
@@
Figure 4: Dependence on the $\eta$-parameter of the excitation energy, calculated on the H4 molecule, in the linear (a) and rectangular (b) geometries, with the 3-21G basis set.
@@
The two approaches CP and LR, at the UHF level of theory can be compared with the exact FCI results. The results are shown in Figs. 3-5, for the molecules H2 and H4 in the linear and rectangular geometry, with the 3-21G basis set.[72]
@@
Figure 5: Dependence on the $\eta$-parameter of the excitation energy, calculated on the H2O, with the STO-3G basis set. In Fig. 5(b), a zoom on the differences between the different approaches in the region around $\eta=1$ is plotted.
@@
Figure 6: Dependence on the $\eta$-parameter of the excitation energy, calculated on H2O. In Fig. 6(a) the 6-31G basis set is used, while in Fig. 6(b), the $c$-pVDZ basis set is used. At $\eta=0.2$ and $\eta=0.6$, the TD-HF algorithm in PySCF was not able to converge due to high linear dependencies.
@@
For all cases, the LR results are closer to the FCI ones. The plots also show that, in the weakly-interacting limit ($\eta\to 0$), the LR formalism coincides with FCI at first order in $\eta$ as expected from Eqs. (45) and (51), while the two approaches differ in the interacting case ($\eta=1$). In the water example with the STO-3G (Fig. 5),[71] 6-31G,[74, 75] and cc-pVDZ[76] basis sets (Fig. 6), the three methods show small discrepancies across the range $\eta\in[0,1]$.